1/1

Lord George Gordon Byron ,

Don Juan

CANTO THE FIRST

I want a hero: an uncommon want, When every year and month sends forth a new one, Till, after cloying the gazettes with cant, The age discovers he is not the true one; Of such as these I should not care to vaunt, I ’ll therefore take our ancient friend Don Juan— We all have seen him, in the pantomime, Sent to the devil somewhat ere his time.

Vernon, the butcher Cumberland, Wolfe, Hawke, Prince Ferdinand, Granby, Burgoyne, Keppel, Howe, Evil and good, have had their tithe of talk, And fill’d their sign posts then, like Wellesley now; Each in their turn like Banquo’s monarchs stalk, Followers of fame, ’nine farrow’ of that sow: France, too, had Buonaparte and Dumourier Recorded in the Moniteur and Courier.

Barnave, Brissot, Condorcet, Mirabeau, Petion, Clootz, Danton, Marat, La Fayette, Were French, and famous people, as we know: And there were others, scarce forgotten yet, Joubert, Hoche, Marceau, Lannes, Desaix, Moreau, With many of the military set, Exceedingly remarkable at times, But not at all adapted to my rhymes.

Verkinto de tiu ĉi Angla poemo estas Lord George Gordon Byron (∗1788-01-22 – †1824-04-18).

Arg-1019-2062 (2014-05-20 22:30:37)

http://www.poezio.net/version?poem-id=1019&version-id=2062